I do not know how to complete this table and I need some help.
The important info is: A chemist has one solution of citric acid that is 20% acid and another solution of citric acid that is 80% acid. She plans to mix these solutions together to make 200 liters of a solution that is 50% acid.
And the question is: Complete the table to show the liters of 20% and 80% solutions that will be used to make the 50% solution. Use x to represent the number of liters of the 80% solution that will be used to make the 50% solution.
Liters of Solution Liters of Acid
20% Solution
80% Solution x
50% Solution 200 0.50(200)

Answer by mananth(16946)

---------percent ---------------- quantity
Solution I 80 ---------------- x liters
Solution II 20 ------ 100 - x liters
Mixture 50.00% ---------------- 100

80x+20(100-x)=50*100

80x+2000-20x=5000
80x-20x=5000-2000
60x=3000
/60
x=50 liters 80.00% Solution I
50 liters 20.00% Solution II